OpenStudy (anonymous):

One card is selected at random from a deck of cards. Determine the probability that the card is a 3 or 7.

OpenStudy (anonymous):

P ( A or B ) = P (A) + P (B) - P ( A and B ) P ( 3 or 7 ) = P (3) + P (7) - P ( 3 and 7 ) = 4/52 + 4/52 - 0 (a card can't be both) = 1/13 + 1/13 = 2/13